OpenStudy (anonymous):

v+9/3=8 two step equation

OpenStudy (anonymous):

v+9 over 3

OpenStudy (anonymous):

5 man

OpenStudy (anonymous):

\[\large{\frac{v + 9}{3} = 8}\]\[\large{\frac{v}{3}+ \frac{9}{3} = 8}\]\[\large{\frac{v}{3}+3 = 8}\]\[\large{\frac{v}{3} = 8 - 3}\]\[\large{\frac{v}{3} = 5}\]\[\large{3\left(\frac{v}{3}\right) = 3(5)}\]\[\boxed{\large{v = 15}}\]

OpenStudy (anonymous):

oh, now I get it. Excuse me for that.

OpenStudy (anonymous):

no problem